OpenStudy (anonymous):

Finding Percentile... 1,2,3,4,5,8,9,15 So im doing a sample problem i made up and i end up with 2.7 after using the percentile formula. However it isnt listed in the data, so i decided to average the 2nd and 3rd values and got 2.5 which also is not in the data.. So im wondering what am i doing wrong?

OpenStudy (anonymous):

Oh and im looking for the 30%

OpenStudy (amistre64):

how many numbers are there?

OpenStudy (anonymous):

8

OpenStudy (amistre64):

then a percentile is 8/100; or simply .08 how many of those do we need? 30 right?

OpenStudy (amistre64):

.08(30) = 2.4 we like decimals, they make this easier; round up

OpenStudy (anonymous):

Oh interesting approach. I was using v= 30/100 (n +1)

OpenStudy (amistre64):

ah, a sample ... i considering a population

OpenStudy (amistre64):

same concept tho :) it looks like we want the (2.xx position) so the 3rd position would be the 30th percentile

OpenStudy (anonymous):

Oh gotchya. Basically everything below the 3rd value falls in the 30th percentile.. Thanks i get it now =)

OpenStudy (amistre64):

good luck :)

OpenStudy (amistre64):

if we didnt have a decimal, the we would have to average some positions

OpenStudy (anonymous):

Like if it was 2.7997979797977979 or whatever?

OpenStudy (amistre64):

lol, no say we had 100 numbers (just to make this simple) 100/100 = 1, 30% is 30 the 30th position needs to be included inside of the 30th percentile, so we would have to interpolate between the 30th and 31st position values

OpenStudy (amistre64):

ideally that would be (30th + 31st)/2, but different texts like different exactnesses

OpenStudy (anonymous):

I see what your saying but i feel like i would have trouble distinguishing when to average or when to round up.

OpenStudy (amistre64):

if we have a decimal expansion ... we round to the next position since there is no say: 2.34675 th position if our nth percentile number is an interger, we need to average the nth and (n+1)th positions

OpenStudy (amistre64):

its like finding the median; the number in the middle is either in the middle, or the average of the 2 on the side of the middle

OpenStudy (anonymous):

Ok. So if i got 2 id need to find the average between that and the next value. But if i get a decimal i just round up.

OpenStudy (amistre64):

correct as an exercise lets include a "m"edian in your set, and find the 50% percentile of the 9 values 1,2,3,4,m,5,8,9,15 9/100 * 50 = 4.5 ; our median is "m" where we expect it to be lets take it out and find the 50% of 8 values :) 1,2,3,4,5,8,9,15 ^ 8/100 * 50 = 4 , we want to average the 4th and 5th postions

OpenStudy (anonymous):

This totally makes sense. 4.5 Is the m

OpenStudy (anonymous):

Everything below that falls into the 50th percentile

OpenStudy (anonymous):

Thx again

OpenStudy (amistre64):

youre welcome :)